OpenStudy (anonymous):

if carmen uses 1/2 lb of lettuce, 1 1/14 lb of tomatoes, and 1 1/3 lb of cucumber to make a salad. how many pounds of produce does carmen use in all

OpenStudy (anonymous):

(1/2) + (15/14) + (4/3) = (21/42) + (45/42) + (56/42) (by expressing each over a common denominator) = 61/21